ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 114, 1999

Arad, Z., E. Fisman and M. Muzychuk Generalized table algebras  29

Berele, A. and J. R. Stembridge Denominators for the Poincaré series of invariants of small matrices

157

Burke, D. R. and Y. Matsubara The extent of strength in the club filters

253

Calvaruso, G., D. Perrone and L. Vanhecke Homogeneity on three-dimensional contact metric manifolds

301

Campbell, H. E. A. and N. E. Kechagias Arnon completion of the Dyer–Lashof algebra

189

Demir, H. See O. Sh. Mukhtarov

Fisman, E. See Z. Arad

Grunewald, F. and J. Schwermer On the concept of level for subgroups of SL_{2} over arithmetic rings

205

Kechagias, N. E. See H. E. A. Campbell

Lefcourt, T. R. Galois groups and complete domains  323
Levin, A. L. and E. B. Saff Fast decreasing rational functions  125

Manning, A. and K. Schmidt Common homoclinic points of commuting toral automorphisms

289

Matoušek, J. On embedding trees into uniformly convex Banach spaces

221

Matsubara, Y. See D. R. Burke

Mukhtarov, O. Sh. and H. Demir Coerciveness of the discontinuous initial-boundary value problem for parabolic equations

239

Muzychuk, M. See Z. Arad

Navarro, G. and L. Sanus A partition of characters associated to nilpotent subgroups

359

Nicolaescu, L. I. Eta invariants of Dirac operators on circle bundles over Riemann surfaces and virtual dimensions of finite energy Seiberg–Witten moduli spaces

61

Perrone, D. See G. Calvaruso

Ponnusamy, S. and F. Rønning Integral transforms of functions with the derivative in a halfplane

177

Rønning, F. See S. Ponnusamy Rubin, B. Fractional integrals and wavelet transforms associated with

Blaschke–Levy representations on the sphere

1

Saff, E. B. See A. L. Levin Sanus, L. See G. Navarro Schmidt, K. See A. Manning Schwermer, J. See F. Grunewald

Shioya, M. Splitting P_{κ}λ into maximally many stationary sets  347

Stembridge, J. R. See A. Berele Vanhecke, L. See G. Calvaruso Weit, Y. On Furstenberg’s characterization of harmonic functions on

symmetric spaces

265

Yu, J. S. See Z. Zhou Zhou, Z. and J. S. Yu Linearized oscillation theorems for neutral

difference equations

149

Ziegler, T. An application of ergodic theory to a problem in geometric Ramsey theory 271